Current Attempt in Progress Of all the times this hard drive could crash, it had to be now,Betty cried. How can I finish tQuestion 6 of 10 < -71 MIT (1) Calculate the begininng finished goods. (Round answer to the nearest dollar amount, e.g. 5,275(6) Calculate the direct material used. (Round answer to the nearest dollar amount, e.g. 5,275.) Direct material used $ (7) C
Answer #1
Details We have: Particulars Amt in$
1 Total Inventory at the end of May 1,00,000.00
2 FG inventory 80% of Total Inventory May      80,000.00
3 Closing DM Inventory June      40,000.00
4 Closing WIP Inventory June      21,000.00
5 Closing FG Inventory June      45,000.00
6 Increase in DM Inventory June by 25% from May 25%
7 Labour Cost paid During June      23,000.00
8 Sales Turnover During June 2,91,000.00
9 OH rate is 200% of DL Cost 200%
10 Product Pricing at 100% mark up COGS 100%
Answers:
1 Beginning of Finished Goods      80,000.00
80% of Total Inventory of May end of $ 100,000/-
2 Beginning of Direct Raw Materials      32,000.00
June DM Inventory Increased by 25% of May End DM Inventory
3 Beginning of Work In Process    -12,000.00
Loss of Inventory
4 Cost of Goods Sold 1,45,500.00
100% mark up on COGS for Sell Price, Total Turnover $ 291,000/-
5 Cost of Goods Manufactured June      99,500.00
Excluding Overhead cost which is 200% of DL Cost, which comes to (23000*200%)= 46,000
6 Direct Materials Used      76,500.00
Excluding Direct Labour cost during June 23000, balance is Direct Materials Used COGM
7 Amount of Purchases      84,500.00
Opening DM June 32000 a
Amount of Purchases During Month (Balancing Figure) 84500 b
Total of Above 116500 (a+b)=c
Material Used During June 76500 d
Balance of DM at end of June 40000 (c-d)=e
8 Direct Labour for the Month      23,000.00
Labour cost is incurred during month at the same time
9 Amount of Applied Overhead      46,000.00